# Enhancing predictive capabilities in fusion burning plasmas through surrogate-based optimization in core transport solvers

### 19 Dec 2023 v1

@@ 3.3 Restarting simulations for parameter scans
This technique was exploited for the study of the parameter space of the SPARC tokamak Creely2020a ; Rodriguez-Fernandez2022 when operating in L-mode-like conditions. We utilized nonlinear CGYRO Candy2016 to evaluate core transport at 6 radial locations ($r/a=0.35,0.55,0.75,0.875,0.9,0.943$) and scanned auxiliary input power, edge density boundary condition and edge temperature boundary condition. As shown in Fig. 7, first convergence took 12 profile evaluations, but subsequent scenarios required an average of 5 extra evaluations for convergence in the three predicted channels ($T_{e}$, $T_{i}$, $n_{e}$). This unique study enabled the investigation of 12 scenarios at a total cost of 71 profile evaluations (426 local $\delta f$ CGYRO simulations). Physics results from this and other applications of PORTALS will be part of upcoming publications and conference presentations.
@@ 4 Guidelines for accurate profile predictions
At the time of publication, we have utilized this workflow for the nonlinear gyrokinetic profile prediction of 29 different plasma conditions, many of which were for burning plasmas. This section is intended to outline “best practices” that have been determined from these efforts.
@@ 4.3 Ion-scale simulations
Validation studies in current experiments, with widely varying heating schemes, collisionality regimes and plasma parameters, require additional work to understand the validity of ion-scale gyrokinetics to reproduce core transport. Performing linear stability and numerical resolution scans around experimentally-measured conditions is an approach that is often taken in the community (e.g. Ref. howard_2024_pop for predictions of DIII-D plasmas).
@@ 4.4 Ion bundling
Similarly to the impurity lumping, it is observed howard_2024_pop that the effect of separating hydrogenic ions into deuterium and tritium is marginal on core turbulence modeling of burning plasmas and therefore it becomes advantageous to simulate a lumped $A=2.5$ hydrogenic main ion. Studies of isotope effect in current machines Belli2021 show that using an effective main ion mass can reproduce many aspects of the turbulence. However, the effect on alpha heating and fusion production is, of course, retained during the calculation of targets.
@@
To demonstrate the efficiency of PORTALS, here we aim at simultaneously bringing 5 channels (electron temperature, ion temperature, electron density, lithium impurity density and angular rotation) to steady-state of the DIII-D ITER Similar Shape (ISS) H-mode plasma from Ref. howard_2024_pop , with ion-scale, nonlinear CGYRO Candy2016 and NEO Belli2008 simulations providing the turbulent and neoclassical components of cross-field transport. We allow radiation (Bremsstrahlung, line and synchrotron) and energy exchange (classical and turbulent) to evolve. We employ the 5 radial locations described in Sec. 4.1 ($r/a=[0.35,0.55,0.75,0.875,0.9]$, as illustrated in Fig. 1 for this specific DIII-D ISS case) but calculate targets in a higher resolution grid of 20 points from $r/a=0.0$ to $0.9$.
Nonlinear gyrokinetic simulations using the CGYRO code were performed at each of the radial locations. The general simulation setup is similar to that reported in Ref. howard_2024_pop with some modest changes that are meant to lead to improve the turbulence statistics and therefore lead to more reliable time averaged heat and particle fluxes. Although the exact box sizes varied slightly across the radius, the nominal simulation setup was targeting box sizes in the radial and bi-normal direction of $\sim 120\times 120\rho_{s}$ for $L_{x}$ and $L_{y}$. This simulation domain was represented with 24 toroidal modes ($n_{n}$) spanning from $0.053$ to $1.219$ in $k_{y}\rho_{s}$. Nominally 512 radial modes ($n_{r}$), 24 toroidal modes ($n_{n}$), 24 theta points ($n_{\theta}$), 24 pitch angles $(n_{\xi})$ and 8-12 energies ($n_{energy}$) were used, with the exact number of radial modes varying somewhat depending on the radial location simulated. Simulations utilized high accuracy collisions implemented with the Sugama collision operator Sugama1997 , realistic geometry implemented with the Miller Extended Harmonic formulation Arbon2020 , electromagnetic turbulence ($\delta\phi$, $\delta A_{\|}$, $\delta B_{\|}$), and included 5 gyrokinetic species (D, T, Lumped impurity species, trace Li). The addition of the trace Li species ($n_{Li}/n_{e}=1\cdot 10^{-6}$) is to allow for the prediction of the lithium density profiles in the simulation.
@@
Fig. 12 shows the evolution of the 5 kinetic profiles from the original parameterized profiles (red) howard_2024_pop to the multi-channel flux-matched ones (green). Starting transport fluxes (red lines in bottom subplots) are very far from targets, but after 26 transport evaluations per radius (130 nonlinear CGYRO simulations), the plasma is brought to steady-state in all coupled channels simultaneously. For visualization purposes, transport vs target fluxes are plotted in both Fig. 12 and Fig. 13 with different y-axis range.
@@
Fig. 14 shows the time traces of the CGYRO $\delta f$ simulations at $r/a=0.55$ as representative location. Simulations are restarted from a baseline case (blue shaded region) that was run for $\sim 360a/c_{s}$ and each of them is let to evolve to their corresponding saturated state. At this location it was assumed that running for an additional $\sim 400a/c_{s}$ was sufficient to evaluate the saturated state, which was determined by visual inspection, but metrics for automation of PORTALS are under investigation. The mean flux in the saturated state is determined by the mean of the time trace in the saturated region, and the confidence bounds are $2\sigma_{S}$, where $\sigma_{S}=\frac{\sigma}{\sqrt{N}}$. Here, $\sigma$ is the standard deviation of the time trace and $N$ is the number of independent samples, found by decomposing the full time window into periods of $3\tau_{C}$, where $\tau_{C}$ is the autocorrelation time. This assumes that time points separated by more than 3 autocorrelation periods can be considered to be independent from each other and contribute to the reduction of the total error of the time signal. Here we also make the assumption that target and neoclassical fluxes have no error associated to their calculation, which could be included in future work when constructing better metrics to determine convergence in PORTALS simulations.
@@ 6.1 On the need for self-consistent, multi-channel predictions
Figure 16: Example of predicted lithium particle transport coefficients with CGYRO for the DIII-D simulation presented in Ref. howard_2024_pop . Coefficients calculated with background profiles from (black) $Q_{i}$-matched, standalone simulations and (red) 3-channel profile predictions.
To illustrate this, Fig. 16 displays the differences of the calculated impurity particle transport coefficients with CGYRO between using the self-consistent, 3-channel predicted profiles or only using the local, standalone, $Q_{i}$-matched gradients. This example correspond to the DIII-D ISS plasma howard_2024_pop . Differences are such that can lead incorrect validation conclusions if not self-consistent results are compared to experimentally-inferred transport coefficients.
@@ 6.2.1 Potential for future work
While the PORTALS techniques have shown promising results and great performance in the predictions using nonlinear CGYRO, there are potential avenues for improvement. As part of future work, further development of the surrogate-based optimization techniques will involve a more careful selection and investigation of Bayesian optimization aspects to further increase the efficiency of the formulation. These may include physics-informed GP surrogate models (e.g. that capture critical gradient behavior), other acquisition functions (accounting for the exploration / exploitation tradeoff) and better metrics to determine if converged-gradients solutions have been achieved.
@@ 7 Conclusions and Prospects
This paper has presented the approach of formulating the multi-channel steady-state transport equations in tokamak geometry as a surrogate-based optimization problem. The implementation of this approach in PORTALS is demonstrated to reduce the number of $\delta f$ transport simulations required to bring the plasma to flux-matching conditions, as compared to standard methods. Advantages of the framework were discussed, particularly those relevant to power plant design and burning plasma simulations.
@@ Acknowledgments
The authors would like to thank W. Dorland, N.R. Mandell and A. di Siena for fruitful conversations on high-fidelity transport solvers. We also appreciate discussions with colleagues during the 31st and 32nd US Transport Taskforce meetings (2022-2023) and the 2nd Meta Adaptive Experimentation workshop (2023). We thank E. Belli and G. Staebler for the development of CGYRO and TGLF, respectively. This work was funded by Commonwealth Fusion Systems under RPP020 and US DoE under grants DE-SC0017992, DE-SC0014264, DE-SC0024399 and DE-SC0018287. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a US Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02–05CH11231, for the PORTALS-CGYRO simulations. Clusters hosted at the Massachusetts Green High Performance Computing Center (MGHPCC) were used to perform the PORTALS-TGLF benchmark simulations of SPARC and ARC (MIT-NSE and MIT-RPP partitions) and of DIII-D, ITER and JET (MIT-PSFC partition, funded by DoE grant number DE-FG02-91-ER54109). GPT-4 was used to enhance parts of the manuscript for clarity and coherence, and to assist in the creation of Fig. 1a.

### 09 Apr 2024 v2

@@ 3.3 Restarting simulations for parameter scans
This technique was exploited for the study of the parameter space of the SPARC tokamak Creely2020a ; Rodriguez-Fernandez2022 when operating in L-mode-like conditions rodriguez-fernandez_core_2024 . We utilized nonlinear CGYRO Candy2016 to evaluate core transport at 6 radial locations ($r/a=0.35,0.55,0.75,0.875,0.9,0.943$) and scanned auxiliary input power, edge density boundary condition and edge temperature boundary condition. As shown in Fig. 7, first convergence took 12 profile evaluations, but subsequent scenarios required an average of 5 extra evaluations for convergence in the three predicted channels ($T_{e}$, $T_{i}$, $n_{e}$). This unique study enabled the investigation of 12 scenarios at a total cost of 71 profile evaluations (426 local $\delta f$ CGYRO simulations) rodriguez-fernandez_core_2024 . Physics results from this and other applications of PORTALS will be part of upcoming publications and conference presentations.
@@ 4 Guidelines for accurate profile predictions
At the time of publication, we have utilized this workflow for the nonlinear gyrokinetic profile prediction of over 30 different plasma conditions, many of which were for burning plasmas. PORTALS has been employed to perform direct nonlinear gyrokinetic simulations with CGYRO Candy2016 to predict steady-state plasmas in SPARC PRD rodriguez-fernandez_nonlinear_2022 , ITER Baseline Scenario howard_2024_iter , DIII-D ITER Similar Shape howard_simultaneous_2024 , ARC holland_2023_aps , SPARC L-modes rodriguez-fernandez_core_2024 and JET H, D and T Ohmic plasmas prf_2023_eps . These studies are part of a broader effort to both validate nonlinear gyrokinetics in current experiments and to predict the performance of future devices. This section is intended to outline “best practices” that have been determined from these efforts.
@@ 4.3 Ion-scale simulations
Validation studies in current experiments, with widely varying heating schemes, collisionality regimes and plasma parameters, require additional work to understand the validity of ion-scale gyrokinetics to reproduce core transport. Performing linear stability and numerical resolution scans around experimentally-measured conditions is an approach that is often taken in the community (e.g. Ref. howard_simultaneous_2024 for predictions of DIII-D plasmas).
@@ 4.4 Ion bundling
Similarly to the impurity lumping, it is observed howard_simultaneous_2024 that the effect of separating hydrogenic ions into deuterium and tritium is marginal on core turbulence modeling of burning plasmas and therefore it becomes advantageous to simulate a lumped $A=2.5$ hydrogenic main ion. Studies of isotope effect in current machines Belli2021 show that using an effective main ion mass can reproduce many aspects of the turbulence. However, the effect on alpha heating and fusion production is, of course, retained during the calculation of targets.
@@
To demonstrate the efficiency of PORTALS, here we aim at simultaneously bringing 5 channels (electron temperature, ion temperature, electron density, lithium impurity density and angular rotation) to steady-state of the DIII-D ITER Similar Shape (ISS) H-mode plasma from Ref. howard_simultaneous_2024 , with ion-scale, nonlinear CGYRO Candy2016 and NEO Belli2008 simulations providing the turbulent and neoclassical components of cross-field transport. We allow radiation (Bremsstrahlung, line and synchrotron) and energy exchange (classical and turbulent) to evolve. We employ the 5 radial locations described in Sec. 4.1 ($r/a=[0.35,0.55,0.75,0.875,0.9]$, as illustrated in Fig. 1 for this specific DIII-D ISS case) but calculate targets in a higher resolution grid of 20 points from $r/a=0.0$ to $0.9$.
Nonlinear gyrokinetic simulations using the CGYRO code were performed at each of the radial locations. The general simulation setup is similar to that reported in Ref. howard_simultaneous_2024 with some modest changes that are meant to lead to improve the turbulence statistics and therefore lead to more reliable time averaged heat and particle fluxes. Although the exact box sizes varied slightly across the radius, the nominal simulation setup was targeting box sizes in the radial and bi-normal direction of $\sim 120\times 120\rho_{s}$ for $L_{x}$ and $L_{y}$. This simulation domain was represented with 24 toroidal modes ($n_{n}$) spanning from $0.053$ to $1.219$ in $k_{y}\rho_{s}$. Nominally 512 radial modes ($n_{r}$), 24 toroidal modes ($n_{n}$), 24 theta points ($n_{\theta}$), 24 pitch angles $(n_{\xi})$ and 8-12 energies ($n_{energy}$) were used, with the exact number of radial modes varying somewhat depending on the radial location simulated. Simulations utilized high accuracy collisions implemented with the Sugama collision operator Sugama1997 , realistic geometry implemented with the Miller Extended Harmonic formulation Arbon2020 , electromagnetic turbulence ($\delta\phi$, $\delta A_{\|}$, $\delta B_{\|}$), and included 5 gyrokinetic species (D, T, Lumped impurity species, trace Li). The addition of the trace Li species ($n_{Li}/n_{e}=1\cdot 10^{-6}$) is to allow for the prediction of the lithium density profiles in the simulation.
@@
Fig. 12 shows the evolution of the 5 kinetic profiles from the original parameterized profiles (red) howard_simultaneous_2024 to the multi-channel flux-matched ones (green). Starting transport fluxes (red lines in bottom subplots) are very far from targets, but after 26 transport evaluations per radius (130 nonlinear CGYRO simulations), the plasma is brought to steady-state in all coupled channels simultaneously. For visualization purposes, transport vs target fluxes are plotted in both Fig. 12 and Fig. 13 with different y-axis range.
@@
Fig. 14 shows the time traces of the CGYRO $\delta f$ simulations at $r/a=0.55$ as representative location. Simulations are restarted from a baseline case (blue shaded region) that was run for $\sim 360a/c_{s}$ and each of them is let to evolve to their corresponding saturated state. At this location it was assumed that running for an additional $\sim 400a/c_{s}$ was sufficient to evaluate the saturated state, which was determined by visual inspection, but metrics for automation of PORTALS are under investigation. We must note that this approach —given the limited time averaging caused by the otherwise prohibitely cost of running for longer time periods— would not capture changes in the transport fluxes from turbulence and zonal flow patterns that may developed on long time scales, as observed in some gyrokinetic studies weikl_ion_2017 ; rath_transport_2022 . This is generally a limitation of flux-matching frameworks of $\delta f$ simulations that aim to achieve steady-state with high-fidelity gyrokinetic simulations, under the constraint of computational cost. However, most long-time patterns, such as those in the aforementioned papers, typically show early signs (e.g., intermittency or bursts) that might trigger a collapse. These are features that are actively looked for in the simulations used to bring the plasma to steady-state in our workflow, and have not been observed in the specific regimes that PORTALS has been run on so far, including the predictions showed in this paper.
The mean flux in the saturated state is determined by the mean of the time trace in the saturated region, and the confidence bounds are $2\sigma_{S}$, where $\sigma_{S}=\frac{\sigma}{\sqrt{N}}$. Here, $\sigma$ is the standard deviation of the time trace and $N$ is the number of independent samples, found by decomposing the full time window into periods of $3\tau_{C}$, where $\tau_{C}$ is the autocorrelation time. This assumes that time points separated by more than 3 autocorrelation periods can be considered to be independent from each other and contribute to the reduction of the total error of the time signal. Here we also make the assumption that target and neoclassical fluxes have no error associated to their calculation, which could be included in future work when constructing better metrics to determine convergence in PORTALS simulations. This utilization of the estimation of uncertainties in observed data to build more robust surrogate models enabled by the built-in capabilities of the Gaussian processes is one of the key advantages of PORTALS, as discussed in Section 6. The details of the implementation of fixed Gaussian noise (coming from this estimation of the error in the time evolution of the saturated state in the initial value simulations) is further discussed in Ref. rodriguez-fernandez_nonlinear_2022 .
@@ 6.1 On the need for self-consistent, multi-channel predictions
Figure 16: Example of predicted lithium particle transport coefficients with CGYRO for the DIII-D simulation presented in Ref. howard_simultaneous_2024 . Coefficients calculated with background profiles from (black) $Q_{i}$-matched, standalone simulations and (red) 3-channel profile predictions.
To illustrate this, Fig. 16 displays the differences of the calculated impurity particle transport coefficients with CGYRO between using the self-consistent, 3-channel predicted profiles or only using the local, standalone, $Q_{i}$-matched gradients. This example correspond to the DIII-D ISS plasma howard_simultaneous_2024 . Differences are such that can lead incorrect validation conclusions if not self-consistent results are compared to experimentally-inferred transport coefficients.
This work, particularly the capability of PORTALS to bring to steady-state multi-channel profiles, including rotation, opens avenues for more comprehensive validation studies in current experiments. This can be important to, for example, complement validation studies of turbulent transport models to capture intrinsic rotation and momentum transport dynamics grierson_main-ion_2017 ; hornsby_global_2018 ; hornsby_effect_2017 .
@@ 6.2.1 Potential for future work
While the PORTALS techniques have shown promising results and great performance in the predictions using nonlinear CGYRO, there are potential avenues for improvement. As part of future work, further development of the surrogate-based optimization techniques will involve a more careful selection and investigation of Bayesian optimization aspects to further increase the efficiency of the formulation. These may include physics-informed GP surrogate models (e.g. that capture critical gradient behavior and that considers the expected smoothness of the flux responses), other acquisition functions (accounting for the exploration / exploitation tradeoff) and better metrics to determine if converged-gradients solutions have been achieved.
@@
An interesting potential avenue for future work is the exploitation of the uncertainty quantification properties of the GP surrogate models for sensitivity studies of modeling assumptions. The propagation of uncertainties in the assumed models or boundary conditions could provide valuable insights into the robustness of the predictions and their validation against experimental data. Once trained, the GP models can be used to perform sensitivity studies of the input parameters, by finding flux-matched conditions not only for the mean of the posterior distribution, but also for the upper and lower bounds of the confidence intervals. This capability has not been explored yet but will be subject of future investigation.
@@ 7 Conclusions and Prospects
This paper has presented the approach of formulating the multi-channel steady-state transport equations in tokamak geometry as a surrogate-based optimization problem. The implementation of this approach in PORTALS is demonstrated to reduce the number of $\delta f$ transport simulations required to bring the plasma to flux-matching conditions, as compared to standard methods. Advantages of the framework were discussed, particularly those relevant to power plant design and burning plasma simulations. This paper has not delved into any particular transport model validation study nor has provided practical burning plasma predictions —which are done as part of separate publications (e.g. rodriguez-fernandez_nonlinear_2022 ; howard_2024_iter ; howard_simultaneous_2024 ; holland_2023_aps ; rodriguez-fernandez_core_2024 ; prf_2023_eps )—. In describing this unique formulation and solution to the steady-state transport problem, we aimed to offer a comprehensive perspective on PORTALS’ potential and its implications for future fusion research.
@@ Acknowledgments
The PORTALS framework has recently been published open source as part of the MIT Integrated Modeling (MITIM) repository mitim . The authors would like to thank W. Dorland, N.R. Mandell and A. di Siena for fruitful conversations on high-fidelity transport solvers. We also appreciate discussions with colleagues during the 31st and 32nd US Transport Taskforce meetings (2022-2023) and the 2nd Meta Adaptive Experimentation workshop (2023). We thank E. Belli and G. Staebler for the development of CGYRO and TGLF, respectively. This work was funded by Commonwealth Fusion Systems under RPP020 and US DoE under grants DE-SC0017992, DE-SC0014264, DE-SC0024399 and DE-SC0018287. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a US Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02–05CH11231, for the PORTALS-CGYRO simulations. Clusters hosted at the Massachusetts Green High Performance Computing Center (MGHPCC) were used to perform the PORTALS-TGLF benchmark simulations of SPARC and ARC (MIT-NSE and MIT-RPP partitions) and of DIII-D, ITER and JET (MIT-PSFC partition, funded by DoE grant number DE-FG02-91-ER54109). GPT-4 was used to enhance parts of the manuscript for clarity and coherence, and to assist in the creation of Fig. 1a.